# Do MLLMs Capture How Interfaces Guide User Behavior? A Benchmark for Multimodal UI/UX Design Understanding

### v3

@@ Abstract
User interface (UI) design goes beyond visuals, guiding user behavior and overall user experience (UX). Strategically crafted interfaces, for example, can boost sign-ups and drive business sales, underscoring the shift toward UI/UX as a unified design concept. While recent studies have explored UI quality evaluation using Multimodal Large Language Models (MLLMs), they largely focus on surface-level features, overlooking behavior-oriented aspects. To fill this gap, we introduce WiserUI-Bench  [^1] https://github.com/jeochris/wiserui-bench, a novel benchmark for assessing models’ multimodal understanding of UI/UX design. It includes 300 diverse real-world UI image pairs, each consisting of two design variants A/B-tested at scale by actual companies, where one was empirically validated to steer more user actions than the other. Each pair is accompanied one or more of 684 expert-curated rationales that capture key factors behind each winning design’s effectiveness, spanning diverse cognitive dimensions of UX. Our benchmark supports two core tasks: (1) selecting the more effective UI/UX design by predicting the A/B test verified winner and (2) assessing how well a model, given the winner, can explain its effectiveness in alignment with expert reasoning. Experiments across several MLLMs show that current models exhibit limited nuanced reasoning about UI/UX design and its behavioral impact. We believe our work will foster research in UI/UX understanding and enable broader applications such as behavior-aware interface optimization.
@@ 1 Introduction
User interface (UI) design is central to application development. While visual aesthetics lay its foundation, the fundamental goal is to steer user behavior during interaction, such as encouraging sign-ups or purchases, and thereby boost service revenue (Fogg 2002). This behavioral focus expands to UI/UX design, uniting visual form with user experience (UX) considerations (Norman 1988).
@@
Figure 1: A real-world behavior-driven design decision example from our benchmark, grounded in actual A/B test results. Such cases reflect UI changes that steer user actions, making them effective testbeds for evaluating models’ visual reasoning about UI/UX design.
@@
Practitioners validate such decisions through large-scale A/B tests that randomly assign users to alternative UI variants and measure which one more effectively drives the desired action (Kohavi et al. 2009). These tests often reveal that even subtle design changes can significantly lift conversions. For example, as shown in Figure 1, simply reducing the number of options lowered user friction and increased sign-ups. Given that predicting and explaining such outcomes is nontrivial even for humans, hinging on complex cognitive processes of user experience and implicit human preferences, assessing multimodal understanding of UI/UX design becomes a crucial challenge, demanding not only identifications of visual elements but also reasoning about how design influences user behavior.
@@
Benchmark # Samples Unit Source Quality Source Type
Large-Scale User
Validation
Expert
Rationale
Evaluation Objective
Yang and Li (2024)
382 Single Legacy UI Mobile
Guideline Violation
Detection
JitterWeb
(Wu et al. 2024a)
2.3M Pair Synthetic Mobile + Web
Basic Visual Quality
BetterWeb
(Wu et al. 2024a)
892 Pair Synthetic Mobile
Basic Visual Quality
UICrit
(Duan et al. 2024)
983 Single Legacy UI Mobile
UI Design Critique
WiserUI-Bench
300 Pair
Actual
Production
Mobile + Web
(Real A/B Test)
UI/UX Design Understanding
(User Behavior-Oriented)
@@ 2.3M Pair Synthetic Mobile + Web
Recent efforts have explored how to evaluate UI design quality with Multimodal Large Language Models (MLLMs); however, they all fall short in capturing the behavioral impact of UI/UX design, as summarized in Table 1. While some focus solely on basic visual, heuristic-based attributes (Yang and Li 2024; Wu et al. 2024a), others rely on handwritten critiques of individual UIs, lacking validation against real user behavior data and offering limited coverage of cognitive aspects (Duan et al. 2024).
To address these limitations, we introduce WiserUI-Bench (Pairwise UI/UX Design Understanding Benchmark), a novel benchmark for evaluating a model’s multimodal understanding of UI/UX design. WiserUI-Bench includes 300 real-world UI image pairs with winning variants verified by A/B tests from actual companies, spanning varied contexts and sourced from reliable platforms. Each pair is accompanied by one or more expert-curated rationales, 684 in total, written post hoc with knowledge of the winner to explain its effectiveness, covering key behavioral drivers. An example instance is illustrated in the upper part of Figure 2. Uniquely grounded in large-scale real user behavior data, our benchmark mirrors authentic usage and preference patterns, providing a reliable basis for evaluation. Each rationale is tagged with one of the three cognitive UX dimensions articulated by Norman (2007)—perception, what users notice; memory, what they remember; and action, how they act—and with the specific UX law it reflects. This enables fine-grained assessment of how well models grasp the underlying user experience behind each design choice. For example, the action-related rationale in Figure 2 corresponds to Fitts’s Law, which states that users act faster when targets are larger and closer (Fitts and Peterson 1964).
Our benchmark supports two core tasks, illustrated in the lower part of Figure 2: (1) selecting the more effective UI/UX design from a given pair by predicting the A/B test verified winner, and (2) assessing how well model-generated rationales, given the winner, align with expert reasoning. This dual-task setup captures the essential capabilities for UI/UX assessment, not only identifying which design better guides user behavior, but also explaining why reliably.
Experiments on a broad range of proprietary and open-source MLLMs on WiserUI-Bench show that current models struggle with behavior-centered visual reasoning. Models perform near-random on selection, and their post-hoc reasoning remains insufficient to reach expert level. We expect that WiserUI-Bench and our novel UI/UX-focused tasks will spur further research on this critical problem and potentially push toward behavior-aware UI generation, the direction that truly matters for achieving user-centered design.
In summary, our main contributions are as follows:
  1. 1.
@@
    We release WiserUI-Bench, a benchmark of 300 real-world UI pairs uniquely constructed from verified results of actual large-scale A/B tests conducted by real companies, along with expert-curated key rationales.
  3. 3.
@@
    Through comprehensive experiments and analyses across diverse MLLMs, we uncover their limited ability to reason about behavior-driven UI effectiveness, highlighting a critical gap in current visual reasoning.
2 Related Works
Visual Reasoning
@@ 2 Related Works
Visual reasoning benchmarks for Multimodal Large Language Models (MLLMs) evaluate models’ ability to process visual information and reason at a cognitive level. While some target general tasks (Yue et al. 2024; Liu et al. 2024b; Fu et al. 2024; Li et al. 2024a), others focus on specific domains like mathematics or science (Hao et al. 2025; Lu et al. 2024; Xu et al. 2025). However, few benchmarks address reasoning about human preferences in visual contexts. Moreover, although some studies explore visual reasoning across multiple images (Cheng et al. 2025; Zhao et al. 2024), such cases remain rare. Our task incorporates both aspects, reasoning over paired UI images and aligning with behavior-driven preferences, as captured by real-world A/B test outcomes.
Figure 2: Overview of WiserUI-Bench and two main tasks. Each instance contains a UI image pair with verified A/B test winner and expert-curated key rationales explaining it. These span the three cognitive UX dimensions; the example includes all three, though most cover fewer. MLLMs are evaluated on (1) selecting the more effective UI/UX design by predicting the verified winner, and (2) explaining the effectiveness of a given winner, measured by whether the model covers each expert rationale.
UI Evaluation
@@
Recent advances have driven progress in evaluating UI quality with MLLMs. Yang and Li (2024) proposes a benchmark for detecting guideline violations in individual screens, while Wu et al. (2024a) generates synthetic UI pairs via perturbations such as color noise or spacing changes. However, both are limited to heuristic analysis of surface-level appearance. Duan et al. (2024) evaluates single screens using expert-critique annotations, but the benchmark has key limitations: it lacks validation with real-world user behavior and relies on RICO (Deka et al. 2017), an outdated, mobile-only dataset that limits its relevance to today’s diverse, production-level designs. Prior work also suggests that pairwise evaluation yields more reliable signals, particularly for subjective tasks like ours that target behavior-driven preferences captured by real-world A/B tests (Abel et al. 2020; Bawabe et al. 2021). In this context, WiserUI-Bench offers a stronger foundation by grounding design effectiveness in large-scale user behavior data and requiring models to engage in deeper, nuanced reasoning over both visuals and user experience in a pairwise format.
Simulation of Human Behavior
@@
Efforts to simulate human behavior with AI are ongoing. Park et al. (2023) places LLM-driven characters in a sandbox town, producing socially plausible interactions. Lu et al. (2025a) moves closer to real behavior by using actual platform logs to predict next user actions from prior context, yet remains text-only. In the UI domain, Lu et al. (2025b) and Wang et al. (2025a) deploy persona-based agents to browse interfaces and examine how designs may steer user behavior; the former checks realism only via qualitative interviews, while the latter compares against actual A/B tests but in a single scenario. By contrast, our work grounds evaluation on numerous A/B test results from real users and introduces a novel multimodal task: reason over UI variant pairs to predict user behavior, enabling direct measurement of behavioral fidelity.
@@ 3.1 Benchmark construction
As illustrated in Figure 3, our curation process involves multiple stages to ensure high-quality annotations. First, we extract raw UI images and accompanying textual descriptions from our selected platform sources. Next, we refine these images by eliminating any visual markers, such as arrows, circles, or highlights, to preserve a clean, unbiased evaluation of visual reasoning.
Finally, we construct key rationales that explain the winning variant’s superior performance observed in A/B test results, curated by a group of three experienced UI/UX experts aware of the outcomes. Such expert involvement is essential, as interpreting user behavior often requires deep domain expertise. To minimize the risk of overlooking key behavioral drivers, each rationale was collaboratively developed and refined through multiple rounds of expert review to ensure salience and comprehensive coverage. The process was guided by textual descriptions from data sources as well as well-established principles, including the Nielsen Norman 10 Usability Heuristics (Nielsen 1994) and 12 carefully selected UX laws. These were chosen for their ability to explain how UI elements influence user behavior and capture key aspects of UX. Each rationale is thus tagged with the corresponding law. Since user behavior is often shaped by multiple factors, a single instance may include multiple rationales. See Supplementary for more curation details, including the complete list of UX laws used.
Figure 3: Benchmark construction pipeline. (a) Raw A/B test data were collected from reliable platforms, (b) then UI images were refined for clarity. (c) Finally, essential rationales explaining the test outcomes were crafted by combining insights from both UI/UX experts and data sources.
@@ 3.2 Benchmark Analysis
Our pipeline produces a benchmark of 300 real-world UI image pairs, each with a definitive winner identified through A/B test results and annotated with expert-curated key rationales explaining the core design elements that drove user behavior, totalling 684 entries.
Ensuring diversity in our dataset is critical, as effective UI/UX strategies for guiding user behavior vary widely across contexts (Oinas-Kukkonen and Harjumaa 2018). For example, each page type represents a distinct stage in the user journey, often requiring different design strategies. To capture this, WiserUI-Bench spans 11 page types, 11 industry domains, and 2 device types. Common page types include product/service detail pages (29.0%) and landing pages (18.3%). In terms of industry, retail & e-commerce (43.0%) dominates, reflecting its heavy use of A/B testing for conversion optimization, followed by IT & software (13.7%) and travel & hospitality (13.3%). 84% of the UIs are from web platforms and 16% from mobile. Moreover, since our benchmark is based on real-world UIs, it exhibits a wide variety of design changes. We categorize these into 10 high-level types, with a single instance often involve multiple UI changes. In total, we record 490 distinct UI differences across all samples. For example, the UIs shown in Figure 2 reflect changes in layout & structure (repositioning favorite button) and controls & copy (enlarging the ’add to cart’ button). See Supplementary for detailed statistics.
Figure 4: Rationale statistics for WiserUI-Bench: (a) by UX dimensions and laws; (b) by count per instance.
WiserUI-Bench also exhibits wide variation in the types of reasoning used to explain user behavior, as illustrated in Figure 4. Each instance contains an average of 2.28 rationales, with up to 5 in some cases. The UX laws tagged during curation are grouped into three cognitive dimensions of UX—perception, memory, and action—with four laws assigned to each, drawing on the idea from Norman (2007) that user experience can be interpreted through these dimensions. Each rationale is categorized accordingly, enabling fine-grained evaluation of how well models capture the diverse UX factors underlying each design choice. Refer to Supplementary for more details.
@@ 4 Task Description
Proper evaluation of a model’s multimodal understanding of UI/UX design requires testing not only whether the model can identify the more effective design in guiding user behavior, but also whether it can explain why, grounding its reasoning in the visual and contextual cues. To this end, we define two independent tasks, decision and explanation, as illustrated in Figure 2, both essential for demonstrating genuine understanding. Success requires identifying key visual differences between two UIs, as well as inferring the likely behavioral goal from context (e.g., encouraging sign-ups) and how the interface supports that goal.
@@
The first task assesses whether the model can identify which of two UI/UX designs is more effective at influencing user behavior, with the correct answer determined by large-scale A/B testing run by real companies. Although the task is framed as a binary choice, solving it effectively requires nuanced visual reasoning beyond surface-level comparison.
Task 2: UI/UX Design Rationale Alignment
The second task evaluates a model’s reasoning ability in depth. Unlike the first task, it presents the winning UI/UX design upfront and asks the model to explain the factors contributing to its effectiveness in guiding user behavior. This setup mirrors the condition under which expert rationales were made, providing a fair basis for comparison. The model’s explanation is then compared to expert-curated key rationales, and the degree of alignment indicates the quality of its UI/UX reasoning at an expert-level.

### v4

@@ Abstract
User interface (UI) design goes beyond visuals to shape user experience (UX), underscoring the shift toward UI/UX as a unified concept. While recent studies have explored UI evaluation using Multimodal Large Language Models (MLLMs), they largely focus on surface-level features, overlooking how design choices influence user behavior at scale. To fill this gap, we introduce WiserUI-Bench  [^1] https://github.com/jeochris/wiserui-bench , a novel benchmark for multimodal understanding of how UI/UX design affects user behavior, built on 300 real-world UI image pairs from industry A/B tests, with empirically validated winners that induced more user actions. For future design progress in practice, post-hoc understanding of why such winners succeed with mass users is also required; we support this via expert-curated key interpretations for each instance. Experiments across multiple MLLMs on WiserUI-Bench for two main tasks, (1) predicting the more effective UI image between an A/B-tested pair, and (2) explaining it post-hoc in alignment with expert interpretations, show that models exhibit limited understanding of the behavioral impact of UI/UX design. We believe our work will foster research on leveraging MLLMs for visual design in user behavior contexts.
Do MLLMs Capture How Interfaces Guide User Behavior?
A Benchmark for Multimodal UI/UX Design Understanding
Jaehyun Jeon1, Min Soo Kim1, Jang Han Yoon1, Sumin Shim1, Yejin Choi1, Hanbin Kim1, Dae Hyun Kim1thanks: Co-corresponding authors, Youngjae Yu2footnotemark: 1Yonsei University  2Seoul National University {jaehyun.jeon, dhkim16}@yonsei.ac.kryoungjaeyu@snu.ac.kr
@@ 1 Introduction
User interface (UI) design is central to application development. While visual aesthetics lay its foundation, the fundamental goal is to steer user behavior during interaction, such as encouraging sign-ups or purchases, and thereby boost service revenue (Fogg, 2002). This behavioral focus expands to UI/UX design, uniting visual form with user experience (UX) considerations (Norman, 1988).
@@
Figure 1: A real-world, behavior-aware design decision example from WiserUI-Bench, grounded in A/B test results, illustrating how UI changes steer user actions.
@@
Practitioners validate such decisions through large-scale A/B tests that randomly assign users to alternative UI variants and measure which design more effectively drives the desired action (Kohavi et al., 2009). For example, as shown in Figure 1, version B achieves higher sign-ups overall, interpreted as reduced user friction from fewer options. Such interpretation reflects an abductive reasoning process (Hobbs et al., 1993; Bhagavatula et al., 2020), inferring the most plausible explanations for aggregate outcomes that are used to guide subsequent design decisions (Quin et al., 2024). Given that predicting and explaining these results is nontrivial even for humans, hinging on complex cognitive processes underlying user experience, assessing multimodal understanding of UI/UX design becomes a crucial challenge, requiring not only identifying visual elements but also reasoning about how they influence user behavior.
Figure 2: Overview of WiserUI-Bench and two main tasks. Each instance contains a UI image pair with verified A/B test winner and expert-curated key interpretations explaining it. These span the three cognitive UX dimensions; the example includes all three, though most cover fewer. MLLMs are evaluated on (1) selecting the more effective UI/UX design by predicting the verified winner, and (2) explaining the effectiveness of a given winner, measured by whether the model captures each expert interpretation.
Recent efforts have explored how to evaluate UI design quality with Multimodal Large Language Models (MLLMs); however, they all fall short in capturing the behavioral impact of UI/UX design, as summarized in Table 1. While some focus solely on basic visual, heuristic-based attributes (Yang and Li, 2024; Wu et al., 2024), others rely on handwritten critiques of individual UIs, lacking validation against real user data and offering limited coverage of cognitive aspects (Duan et al., 2024).
To address these limitations, we introduce WiserUI-Bench (Pairwise UI/UX Design Understanding Benchmark), a novel benchmark for evaluating a model’s multimodal understanding of how UI/UX design affects user behavior. WiserUI-Bench comprises 300 real-world UI image pairs with winning variants verified by A/B tests from actual companies, spanning varied contexts and sourced from reliable platforms. Each pair includes expert-curated key interpretations, 684 in total, written post-hoc with knowledge of the winner to explain its effectiveness, covering core behavioral drivers. An example instance is illustrated in the upper part of Figure 2. Grounded in large-scale real user behavior data, our benchmark captures authentic usage and preference patterns, providing a reliable basis for evaluation. Each interpretation is tagged with one of the three cognitive UX dimensions articulated by Norman (2007): perception (what users notice), memory (what users remember), and action (how users act).
Our benchmark supports two core tasks, illustrated in the lower part of Figure 2: (1) selecting the more effective UI/UX design from a given pair by predicting the A/B test-verified winner, and (2) interpreting the UI/UX effectiveness of the given winner by assessing alignment between model-generated and expert-curated interpretations. This task setup captures the essential capabilities for UI/UX assessment, not only to identify which design better guides user behavior, but also to reliably explain why in a post-hoc manner.
Experiments on a broad range of proprietary and open-source MLLMs on WiserUI-Bench show that existing models struggle with visual reasoning in user behavior contexts. Models perform near-random on selection, and their post-hoc interpretations remain insufficient to reach expert level. We expect that WiserUI-Bench and our novel UI/UX-focused tasks will spur further research on this critical and underexplored, user-centered area, advancing behavior-aware reasoning about visual design in MLLMs.
@@
Benchmark # Samples Unit Source Quality Source Type Large-Scale User Validation Expert Interpretation Evaluation Objective
Yang and Li (2024) 382 Single Legacy UI Mobile Guideline Violation Detection
BetterWeb (Wu et al., 2024) 892 Pair Synthetic Mobile Basic Visual Quality
UICrit (Duan et al., 2024) 983 Single Legacy UI Mobile UI Design Critique
WiserUI-Bench (Ours) 300 Pair Actual Production Mobile + Web (Real A/B Test) UI/UX Design Understanding (User Behavior-Focused)
@@
In summary, our main contributions are:
@@
  • We release WiserUI-Bench, a benchmark of 300 real-world UI pairs uniquely constructed from verified results of large-scale industry A/B tests, along with 684 carefully curated expert key interpretations.
@@
  • Through comprehensive experiments and analyses across diverse MLLMs, we uncover their limited ability to understand the behavioral impact of UI/UX design, highlighting a critical gap in current visual reasoning.
2 Related Work
2.1 Visual Reasoning
@@
Visual reasoning benchmarks for Multimodal Large Language Models (MLLMs) evaluate models’ ability to process visual information and reason at a cognitive level. While some target general tasks (Yue et al., 2024; Liu et al., 2024b; Fu et al., 2024; Li et al., 2024b), others focus on specific domains like mathematics or science (Hao et al., 2025; Lu et al., 2024; Xu et al., 2025). However, few benchmarks address reasoning about human behavior or preference patterns in visual contexts. Moreover, although some studies explore visual reasoning across multiple images (Cheng et al., 2025; Zhao et al., 2024), such cases remain rare. Our task incorporates both aspects, reasoning over paired UI images and aligning with behavior patterns, as captured by actual A/B test outcomes.
2.2 UI Evaluation
@@
Recent advances have driven progress in evaluating UI quality with MLLMs. Yang and Li (2024) propose a benchmark for detecting guideline violations in individual screens, while Wu et al. (2024) generates synthetic UI pairs via perturbations such as color noise or spacing changes. However, both are limited to heuristic analysis of surface-level appearance. Duan et al. (2024) evaluates single screens using expert-critique annotations, but it lacks validation with real-world user behavior and relies on RICO (Deka et al., 2017), an outdated, mobile-only dataset that limits its relevance to today’s production-level designs. Prior work also suggests that pairwise evaluation yields more reliable signals, particularly for subjective tasks like ours that target behavior patterns captured by real-world A/B tests (Abel et al., 2020; Bawabe et al., 2021). In this context, WiserUI-Bench offers a stronger foundation by grounding design effectiveness in large-scale user behavior data and requiring models to engage in deeper, nuanced reasoning over both visuals and UX in a pairwise format.
2.3 Simulation of Human Behavior
@@
Efforts to simulate human behavior with AI are ongoing. Park et al. (2023) places LLM-driven characters in a sandbox town, producing socially plausible interactions. Lu et al. (2025a) moves closer to real behavior by using actual platform logs to predict next user actions from prior context, yet remains text-only. In the UI domain, Lu et al. (2025b) and Wang et al. (2025a) deploy persona-based agents to browse interfaces and examine how designs may steer user behavior; the former checks realism only via qualitative interviews, while the latter compares against an actual A/B test but only under a single scenario. By contrast, our work grounds evaluation in numerous real-user A/B test results and introduces diverse multimodal tasks that reason over UI variant image pairs to predict and explain user behavior.
@@ 3.1 Benchmark Construction
Figure 3: Distribution of (a) UI change element types and (b) expert interpretations for WiserUI-Bench.
Figure 4: Benchmark construction pipeline. (a) Raw A/B test data were collected from reliable platforms, (b) then UI images were refined for clarity. (c) Finally, key interpretations explaining the test outcomes were curated by UI/UX experts.
@@
As illustrated in Figure 4, our curation process involves multiple stages to ensure high-quality annotations. First, we extract raw UI images and accompanying textual descriptions from our selected platform sources. Next, we refine these images by eliminating any visual markers that were externally added, such as arrows or circles, to preserve a clean, unbiased evaluation of visual reasoning.
Finally, we construct key interpretations explaining the winning variant’s superior performance observed in A/B test results, curated by three experienced UI/UX experts aware of the outcomes. Such expert involvement is essential, as interpreting user behavior often requires deep domain expertise. To identify core behavioral drivers, only interpretations independently identified by at least two experts were retained. This resulted in final 684 (562 with full, 122 with two-way agreement) out of 845 initial candidates. The process was guided by textual descriptions from data sources and well-established principles, the Nielsen Norman 10 Usability Heuristics (Nielsen, 1994). Each was then tagged with the most relevant law from a curated set of 12 UX laws. Since user behavior is often shaped by multiple factors, a single instance may include multiple interpretations. See Appendix A.1 for details.
@@ 3.2 Benchmark Analysis
Our pipeline produces a benchmark of 300 real-world UI image pairs, each with a definitive winner identified through A/B test results and annotated with expert-curated key interpretations of the core design factors driving user behavior, totaling 684 entries. We include detailed statistics in Appx. A.2.
Ensuring contextual diversity is essential, as effective UI/UX strategies vary widely across use cases (Oinas-Kukkonen and Harjumaa, 2018). For example, page types correspond to different stages of the user journey, requiring distinct strategies. To capture this, WiserUI-Bench spans 11 page types, 13 industry domains, and two device types.
[…]
Table 2: Results of the UI/UX design selection task on WiserUI-Bench, shown overall and by UI change element types. Best and second-best model scores for each metric are marked in bold and underlined, respectively. For CA, the best and second-best UI element type performances within each model are highlighted in blue and yellow, respectively. All metrics are reported as percentages (%).
Moreover, we annotate UI differences using element–attribute associations, specifying the UI element being modified and the corresponding attribute change (e.g., a button with a position change), to capture real-world UI variants with concurrent modifications. On average, each instance contains 1.53 element-level and 1.99 attribute-level annotations, spanning 19 element types and 14 attribute types. We further group the element types by their primary functional roles (Figure 3(a)): Interactive Control, which directly elicits intentional user actions; Content & State Display, which primarily conveys information, semantics, or system state; and Container & Layout Structure, which organizes spatial layout and guides attention at a structural level.
WiserUI-Bench also exhibits diversity in the interpretations used to explain user behavior, as shown in Figure 3(b), with an average of 2.28 interpretations per instance. The tagged UX laws are grouped into three cognitive dimensions of UX, perception, memory, and action, following Norman (2007), with four laws assigned to each.
@@ 4 Task Description
Proper evaluation of a model’s multimodal understanding of UI/UX design requires assessing not only its ability to predict which design more effectively guides user behavior, but also whether it can explain why in a post-hoc manner, grounding its reasoning in visual and behavioral cues. To this end, we define two main tasks, selection and interpretation, as illustrated in Figure 2, both essential for demonstrating genuine understanding. Success requires identifying key visual differences between two UI variants, as well as inferring the underlying behavioral goal from context (e.g., promoting sign-ups) and how the interface supports that goal.
@@
The first task assesses whether the model can identify which of two UI images is more effective at influencing user behavior, with the correct answer determined by large-scale A/B testing run by real companies. Although the task is framed as a binary choice, solving it effectively requires nuanced visual reasoning beyond surface-level comparison.
Task 2: UI/UX Design Interpretation
The second task evaluates a model’s post-hoc interpretation ability on the observed results. Unlike the first task, it presents the winning UI/UX design upfront and asks the model to explain the factors contributing to its effectiveness in guiding user behavior. The model’s explanation is then compared to expert-curated key interpretations, with the degree of alignment serving as a measure of the quality of its expert-level interpretive understanding.